# Towards improved property prediction of two-dimensional (2D) materials using many-body Quantum Monte Carlo methods

### v1

@@ VI Data Availability
Data sharing is not applicable to this article as no new data were created or analyzed in this study.
@@ VII Notes
Please note certain equipment, instruments, software, or materials are identified in this paper in order to specify the experimental procedure adequately. Such identification is not intended to imply the recommendation or endorsement of any product or service by NIST, nor is it intended to imply that the materials or equipment identified are necessarily the best available for the purpose.
@@
D.W. acknowledges the National Institute of Standards and Technology for funding and support. J.A., A.B., P.R.C.K, J.T.K., L.M., B.R. and H.S. were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and the Center for Predictive Simulation of Functional Materials. L.M. also received support (excitonic effects in 2D) from U.S. National Science Foundation grant DMR-2316007. Y.K. was supported by the Basic Science Research Program (2018R1D1A1B07042443) through the National Research Foundation of Korea funded by the Ministry of Education. I.S. acknowledges support by APVV-21-0272, VEGA-2/0131/23 and by the H2020 TREX GA 952165 projects. K.S. and F.A.R. was supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. C.A. acknowledges funding from the National Science Foundation (NSF) under grant number NSF DMR-2213398 and U.S. Department of Energy (DOE) under grant number DE-SC0024236.

### v2

@@ VI Data Availability
Data sharing is not applicable to this article because no new data were created or analyzed in this study.
@@ VII Notes
Please note that certain equipment, instruments, software, or materials are identified in this paper to specify the experimental procedure adequately. Such identification is not intended to imply the recommendation or endorsement of any product or service by the National Institute of Standards and Technology, nor is it intended to imply that the materials or equipment identified is necessarily the best available for the purpose.
@@
D.W. acknowledges the National Institute of Standards and Technology for funding and support. J.A., A.B., P.R.C.K, J.T.K., L.M., B.R., and H.S. were supported by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division as part of the Computational Materials Sciences Program and the Center for Predictive Simulation of Functional Materials. L.M. also received support (excitonic effects in 2D) from US National Science Foundation grant DMR-2316007. Y.K. was supported by the Basic Science Research Program (2018R1D1A1B07042443) through the National Research Foundation of Korea funded by the Ministry of Education. I.S. acknowledges support by APVV-21-0272, VEGA 2/0133/25, VEGA 2/0131/23 and by the H2020 TREX GA 952165 projects and Funding by the EU NextGenerationEU through the Recovery and Resilience Plan for Slovakia under the project Nos. 09I02-03-V01-00012 and 09I05-03-V02-00055. K.S. and F.A.R. were supported by the US Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division. C.A. acknowledges funding from the National Science Foundation under grant NSF DMR-2213398 and the US Department of Energy under grant DE-SC0024236.